Fermi National Laboratory
 

Accelerator Update for February 14, 2007 - February 16, 2007

 

Wednesday February 14

The day shift began with the Tevatron (TeV) recovering from the sector A1 quench, with Operations monitoring the Recycler stash (226.0E10) and the Antiproton Source (Pbar) stack (25.6E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the colliding beams.

Operations prepared the D0 collision hall for access at 8:41 AM.

At 9:33 AM, a Linac RF station (LRF3) tripped off. It reset.

A Pbar expert began an hour long lithium lens study at 10:47 AM.

D0 came out to access around 10:56 AM.

Operations put the TeV through a dry squeeze at 11:54 AM.

Operations put the TeV into shot setup at 12:19 PM.

Preaccelerator experts accessed the H- Source at 2:10 PM to replace an extractor tube. This work halted all beam.

Operations established store 5228 at 2:13 PM with an initial luminosity of 231.03E30.

At 2:28 PM, Operations reported that during a Recycler stash the last three transfers didn’t make it into the Recycler. Operators discovered that an Accumulator kicker (A:EKIK) was having problems. Experts investigated.

Preaccelerator experts came out of access. Operations resumed sending beam to Pbar and NuMI at 2:34 PM.

Operations reported at 4:26 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations reported at 4:41 PM that with expert help a Recycler stash was completed without a hitch and Pbar resumed stacking.

Operations reported at 11:58 PM that the Recycler had stashed four times during the evening shift.

Thursday February 15

The midnight shift began with Operations monitoring store 5228 (70.2E30) and the stash (172.9E10), with Pbar stacking (17.4E10) and NuMI taking beam.

At 12:28 AM, the NuMI beam permit tripped off due to bad beam positions.

At 7:54 AM, the NuMI horn tripped off for the first of two times in two minutes. It reset.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

At 8:47 AM, Operations reported that the Neutron Therapy Facility (NTF) was having a power supply problem for its 32˚ bend magnet. An expert investigated.

At 8:51 AM, the NuMI horn tripped off on a stripline ground fault indication. It reset.

NTF reported at 11:17 AM that they had completed treatment on its last patient.

Preaccelerator experts accessed the H- Source at 11:48 AM to replace an ARC power supply that was not working properly. They completed the job in about 15 minutes.

TeV experts began an end-of-store crystal collimator study at 12:01 PM.

At 1:24 PM, the Booster West Anode Power Supply tripped off. It would not reset remotely. Operators investigated and had it back online within 15 minutes.

A MI expert began a 30-minute study at 1:57 PM.

Operations terminated store 5228 at 2:03 PM.

Cryo System technicians began repairs on a TeV sector A1 wet engine at 2:13 PM. They completed their work in about 30 minutes.

Operations put the TeV into shot setup at 2:38 PM.

At 2:38 PM, the MI beam permit tripped for the first of four times in 11 minutes on bad beam positions. It reset. Operations contacted an expert.

At 2:59 PM, a MI expert pointed out a problem with the RF sector in a MI vacuum program. Operations contacted an expert.

At 3:39 PM, Operations reported a MI RF cavity LCW temperature trip. A MI expert who was already at MI-60 believed it was due to a faulty readback device and bypassed it. He will monitor the system.

Recycler experts began a 20-minute study at 4:39 PM.

MI experts at MI-60 replaced the faulty MI readback module at 4:45 PM.

Operations established store 5229 at 4:50 PM with an initial luminosity of 252.53E30.

At 5:54 PM, Operations reported very slow ACNET response times. Operators investigated and found a problem with a server. Much later it was determined that the electron cooling BPMs were blasting the network due to a bug in a recently updated program. An expert backed out of the program and everything cleared up.

At 6:39 PM, both NuMI and Pbar tripped off. Due to the ACNET problems, it took Operations about 10 minutes before everything was reset.

At 7:06 PM, a NuMI expert asked to have their beam turned off until the readback problem was fixed. This was resolved and operations resumed sending them beam about ten minutes later.

At 7:43 PM, the NuMI beam permit tripped off. It reset.

Operations reported at 11:57 PM that the Recycler had stashed three times during the evening shift.

Friday February 16

The midnight shift began with Operations monitoring store 5229 (103.8E30) and the stash (125.3E10), with Pbar stacking (28.8E10), and with NuMI taking beam.

At 1:46 AM, a TeV quadrupole power supply (T:SQA0) tripped off on a quench indication. Store 5229 took a luminosity hit, but remained intact. Operations notified the experiments. Operators reset the supply and turned it up slowly. It resumed normal running.

At 3:03 AM, CHL reported that the TeV sector A1 cryo system was going bad. Cryo system techs were contacting an expert.

At 3:54 AM, the T:SQA0 power supply tripped off. Operators reset the supply.

At 4:01 AM, T:VA12 tripped off on a ground fault indication. Operations contacted an expert and then investigated. They found a LCW leak in a regulator (HA13) that was located in a rack above these other supplies that had been tripping off. Later, operators pulled out all the corrector power supplies and started drying them with fans.

At 5:29 AM, the TeV quenched at sector F4.

At 6:11 AM, the Pelletron tripped off due to a vacuum pump trip.

At 6:40 AM, a Pbar Debuncher RF station (DRF1) tripped off and wouldn’t reset. Operators made some adjustments and had the station up and running in about 12 minutes.

Operations prepared D0 for access at 7:09 AM.

Operations reported at 8:13 AM that the Recycler stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today are to wait for the TeV cryo to recover and then go into shot setup around noon. The reset of the day and the weekend will be stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/16/2007    email Fermilab